OpenStudy (anonymous):

evaluate the expression for x = 4 and y = 5 6(x^2 - y) -y^2

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

6*(16-5)-25=96-30-25=???????

OpenStudy (alexandervonhumboldt2):

had you fanned me?

OpenStudy (anonymous):

yeah lol

OpenStudy (alexandervonhumboldt2):

thx and so what you got?

OpenStudy (anonymous):

41

OpenStudy (alexandervonhumboldt2):

yeah correct